# Jamming is a first-order transition with quenched disorder in amorphous materials sheared by cyclic quasistatic deformations

### v1

@@ Introduction.
To demonstrate the instability of fragile states, two tests are carried out. First, we perform a compression-decompression perturbation, $\varphi\rightarrow\varphi+\delta\varphi\rightarrow\varphi$, where $\delta\varphi=10^{-8}$, and the energy is minimized after each step. All fragile states ($43<Z<$ 4) become unjammed ($Z=0$) after this perturbation, while jammed states remain. In Fig. 3c, we plot the distribution $p^{\prime}(Z)$ after the compression-decompression perturbation. Independently, without any perturbation we simply regard all fragile states ($43<Z<$ 4) as unjammed ($Z=0$) and recalculate the distribution $\tilde{p}(Z)$. Figure 3c shows that the two distributions $p^{\prime}(Z)$ and $\tilde{p}(Z)$ perfectly coincide.
In the second test, we repeat CAQS simulations by systematically varying the threshold $f_{\rm th}$ in the criterion of the energy minimization algorithm. The fraction $F_{\rm F}$ of fragile states grows and $F_{\rm U}$ of unjammed states decays with increasing $f_{\rm th}$ (Fig. 3d), suggesting that many unjammed states become fragile under a looser force-balance condition $\overline{f}_{i}<f_{\rm th}$ with a larger $f_{\rm th}$. In contrast, the constant $F_{\rm J}$ shows that the definition of jammed states is insensitive to the algorithm parameter. Note that in granular experiments Bi et al. (2011); Zhao et al. (2019), the friction may play the role of $f_{\rm th}$, i.e., the net inter-particle force could be balanced by the frictional force between particles and the supporting plate. According to this assumption, the probability of observing fragile states in experiments would depend on the particle-plate friction that can be changed by replacing the materials. It provides a protocol to examine our scenario of fragile states in future experimental studies.
@@
Figure 3: Stability tests. Average single-particle (a) potential energy $\overline{e}_{i}$ and (b) net force $\overline{f}_{i}$ as functions of $Z$. Lines in (a) and (b) are guides to the eye. (c) Comparison between $p^{\prime}(Z)$ obtained after the compression-decompression perturbation and $\tilde{p}(Z)$ with fragile states counted as unjammed states, for $N=1000$ at $\varphi=0.844$. (d) Fractions of four states as functions of $f_{\rm th}$. Data are obtained for $N=256$ systems, except for (c).
Figure 4: Distribution of the coordination number. (a) Probability distribution $p(Z)$ at a few different $\varphi$ for $N=1000$. (b) $p(Z)$ at $\varphi=0.843$ for a few different $N$. For better visualization, we do not show the unjammed delta-peak at $Z=0$. (c) The fraction of jammed states $F_{\rm J}$ as a function of $\varphi$ for a few different $N$. The intersection of curves gives $\varphi_{\rm J}^{\infty}=0.8432$. (d) The data points of $F_{\rm J}$ with different $N$ collapse as a function of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$. The solid line represents fitting to Eq. (4), with two fitting parameters $u=0.041$ and $\sigma_{\varphi}=0.043$. The average coordination number $\langle Z\rangle$ is plotted as a function of $\varphi$ in (e) and of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$ in (f). The solid line in (f) represents $\langle Z\rangle=F_{\rm J}Z_{\rm J}^{*}$ using $u=0.041$, $\sigma_{\varphi}=0.043$ and $Z_{\rm J}^{*}=4.1$. Symbols in (b-f) have the same meanings.
Scaling analysis of $p(Z)$ near the jamming transition.
The new distribution $p(Z)$ contains two peaks of unjammed and jammed states, whose fractions satisfy $F_{\rm U}(\varphi)+F_{\rm J}(\varphi)=1$ at any $\varphi$. In Figs 4a and b, we plot $p(Z)$ for a few different $\varphi$ and $N$, showing that $p(Z)$ is always double-peaked across the jamming transition, which is typical behavior of a first-order rather than a second-order transition.
To analyze the scaling behavior of $p(Z)$, we consider a general first-order form of first-order phase transitions Binder (1987a, b),
@@
where $\delta(Z)$ and $p_{\rm J}(Z)$ correspond to the unjammed and jammed peaks respectively. The asymptotic jamming density $\varphi_{\rm J}^{\rm\infty}=0.8432$ for $N\rightarrow\infty$ is determined by the intersection of $F_{\rm J}(\varphi)$ curves with different $N$ (see Fig. 4c), close to the J-point density $\varphi_{\rm J}\approx 0.842-0.843$ reported in previous studies O’hern et al. (2003); Vagberg et al. (2011). The difference between $\varphi_{\rm J}^{\rm\infty}=0.8432$ by $\gamma_{\rm max}=0.7$ and $\varphi_{\rm J}^{\rm\infty}=0.8435$ by $\gamma_{\rm max}=1$ (see Extended Data Fig. E1) is too small to conclude any systematic dependence on $\gamma_{\rm max}$. In Ref. Das et al. (2020), an unjamming pocket is reported in the phase diagram for $\gamma_{\rm max}<0.17$, while for larger $\gamma_{\rm max}$, $\varphi_{\rm J}$ is independent of $\gamma_{\rm max}$, consistent with our observations. These results also suggest that the jamming density obtained by cyclic shear is the lowest jamming density on the J-line of all possible jammed states Parisi and Zamponi (2010); Chaudhuri et al. (2010); Babu et al. (2021).
@@
We assume the fraction of jammed states $F_{\rm J}$ having the following scaling form, $F_{\rm J}(\varphi,N)=\mathcal{F}\left[(\varphi-\varphi_{\rm J}^{\infty})N^{\lambda}\right]$. The value of the exponent $\lambda$ is important for understanding the nature of the transition. If the system were thermal, the fraction would be determined by the Boltzmann distribution, $F_{\rm J}\sim\exp\left(\frac{N\delta f}{k_{\rm B}T}\right)$, where $\delta f$ is the single-particle free energy difference between two phases Binder (1987b, a). Because the free energy is non-singular around a first-order phase transition, it can be expanded, giving $\delta f\sim(T-T_{\rm c})$ to the lowest order. Thus, if the jamming transition were a standard first-order phase transition, then $F_{\rm J}(\varphi,N)=\mathcal{F}_{\rm J}\left[(\varphi-\varphi_{\rm J}^{\infty})N\right]$, i.e., $\lambda=1$ (note that $\varphi$ in the granular system is analogous to the temperature $T$ in thermal systems). However, our numerical data can not be collapsed using $\lambda=1$; in contrast, they can be perfectly collapsed using $\lambda=1/2$ (see Fig. 4d).
To understand the origin of $\lambda=1/2$ scaling, recall that the athermal jamming transition is not driven by the free energy difference between the two phases. Below we explain the scaling by the scenario of a first-order transition with quenched disorder. For a finite $N$, due to the presence of disorder in the packing structure, the jamming density $\varphi_{\rm J}^{N}$ should fluctuate around the asymptotic value $\varphi_{\rm J}^{\infty}$. Let us assume a simple Gaussian form of the distribution, $\rho(\varphi_{\rm J}^{N})\sim\exp\left[-\frac{(\delta\hat{\varphi}_{\rm J}+u)^{2}}{2\sigma_{\varphi}^{2}}\right],$ where $\delta\hat{\varphi}_{\rm J}=(\varphi_{\rm J}^{N}-\varphi_{\rm J}^{\infty})N^{1/2}$ follows the standard central limit theorem. Note that $\rho(\varphi_{\rm J}^{N})$ has been explicitly measured in the compression protocol O’Hern et al. (2002); O’hern et al. (2003): the width of $\rho(\varphi_{\rm J}^{N})$ scales as $w\sim N^{-0.55}$ in both 2D and 3D, supporting our assumption. The above assumption also predicts, $\varphi_{\rm J}^{\infty}-\varphi_{\rm J}^{N}\sim N^{-1/2}$, independent of dimensionality. The jammed states are defined by $\varphi>\varphi_{\rm J}^{N}$, and thus
@@
where $\delta\hat{\varphi}=(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$. Equation (4) agrees well with the data in Fig. 4d.
Figure 5: Susceptibilities. We plot (a) the disconnected susceptibility $\chi_{\rm dis}$ rescaled by $N$, (b) the connected susceptibility $\chi_{\rm con}$ rescaled by $N^{1/2}$, and (c) the ratio between $\chi_{\rm dis}$ and $\chi_{\rm con}^{2}$, as functions of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$. The solid black lines in (a-c) are Eqs. (5), (6) and (7) respectively. The dashed red line in (c) is the constant term in Eq. (7), i.e., $\chi_{\rm dis}/\chi_{\rm con}^{2}=\pi\sigma_{\varphi}^{2}/2$. To draw these lines, we use $u=0.036$ and $\sigma_{\varphi}=0.043$ obtained from the fitting in Fig. 4d, and $Z_{\rm J}^{*}=4.1$ determined at $\varphi_{\rm J}^{\infty}=0.8432$ in Fig. 2d. No fitting is performed here. Symbols in (a-c) have the same meanings as indicated in the legend of (a). In (d) and (e), the susceptibilities of jammed states $\chi_{\rm J}$ are plotted as functions of $N$ for a few different $\varphi$, and are fitted to Eq. (8). The open squares in (d) are $\chi_{\rm J}$ data obtained in a small pressure window $0.00110.0009<P<$ 0.0011, corresponding to $\varphi\approx 0.843$. The fitting parameter $\mu$ is plotted in (f) as a function of $\varphi$, where the error bar represents the fitting error.
[…]
@@
The alternative connected susceptibility is defined as $\chi_{\rm con}=d\langle Z\rangle/d\varphi$. From Eq. (3), one finds that $\langle Z\rangle\approx F_{\rm J}Z_{\rm J}^{*}$ and thus $\langle Z\rangle$ has the same scaling form as $F_{\rm J}$, i.e., $\langle Z\rangle(\varphi,N)=\mathcal{Z}[(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}]$, consistent with the data in Fig. 4e and f. Using $\chi_{\rm con}=Z_{\rm J}^{*}\frac{dF_{\rm J}}{d\varphi}$ and Eq. (4), we derive,
@@
To the lowest order, Eq. (7) gives a scaling, $\chi_{\rm dis}\sim\chi_{\rm con}^{2},$ which is a key signature of the presence of quenched disorder. This square relationship appears generally at first-order transitions in disordered systems, such as the first-order phase transition in the RFIM Nattermann (1998), the brittle yielding of amorphous solids Ozawa et al. (2018) and the melting of ultrastable hard-sphere glasses Zhang et al. (2022). Comparing to the simulation data, Eq. (7) works well around the extreme point (see Fig. 5c).
It seems that the non-Gaussian effect in the distribution $\rho(\varphi_{\rm J}^{N})$, which is neglected in the present analysis, is amplified in the data of $\chi_{\rm dis}/\chi_{\rm con}^{2}$, resulting in slight asymmetry. Extended Data Figs. E1-E4 show that the scaling functions Eqs. (4)-(7) work in 3D, and for a different max strain $\gamma_{\max}=1$ in 2D.
Finally let us discuss the finite-size effects of the jammed peak $p_{\rm J}(Z)$ in Eq. (3). For standard first-order phase transitions, $\eta=1/2$ Binder (1987b, a). Thus in that scenario the susceptibility of the jamming peak would scale as, $\chi_{\rm J}\equiv N\sigma_{\rm J}^{2}=N\langle(Z-Z_{\rm J}^{*})^{2}\rangle_{\rm J}\sim N^{0}$, where $\sigma_{\rm J}^{2}$ is the variance of $p_{\rm J}(Z)$ and $\langle\ldots\rangle_{\rm J}$ represents the average restricted to the jammed states only. However, our $\chi_{\rm J}$ data disagree with this scaling (see Figs.  5d-f). At different $\varphi$, $\chi_{\rm J}$ can be fitted to a pow-law form,
@@
where $\mu=d(2\eta-1)$ is a fitting parameter. At large or small $\varphi$ away from $\varphi_{\rm J}^{\rm\infty}$ (see Fig.  5f), the exponent is close to zero ($\mu\approx-0.2$), suggesting that the local coordination numbers $Z_{i}$ are uniformly distributed. The deviation from the uncorrelated behavior ($\mu=0$) is significant around $\varphi_{\rm J}^{\rm\infty}$, where $\mu\approx-0.8$ reaches the minimum.
@@
The result of $\mu$ shall be interpreted with care. In general, (i) $\mu=0$ corresponds to uniformity of $Z_{i}$, (ii) $\mu>0$ to hyperuniformity with a vanishing $\chi_{\rm J}$ in the thermodynamic limit, and (iii) $\mu<0$ to hyperfluctuations with a diverging $\chi_{\rm J}$ in the thermodynamic limit that typically appears at the critical point in a second-order phase transition. However, the negative $\mu$ in Fig. 5f is not due to the criticality of the jamming transition. Here it is essential to consider the volume fluctuations. The negative $\mu$ is obtained from a $\varphi$-controlled setup in our CAQS simulations. If we instead select configurations around a constant pressure $P$, then the fluctuations become significantly smaller, and $\mu\approx 0$ (Fig. 5d). This observation is consistent with a previous study Ikeda (2023), suggesting that the large fluctuation of $Z$ near $\varphi_{\rm J}^{\rm\infty}$ in the $\varphi$-controlled protocol might be originated from the fluctuation of $\varphi_{\rm J}^{N}$. In fact, careful measurement gives $\mu\simeq 1$ in a $P$-controlled protocol at jamming, confirming hyperuniformity Hexner et al. (2018, 2019). We emphasize that the finite-size effects of the jammed peak $p_{\rm J}(Z)$ contribute negligibly to the scaling of $F(\varphi,N)$, $\chi_{\rm dis}(\varphi,N)$ and $\chi_{\rm con}(\varphi,N)$. In other words, the first-order nature of the jamming transition is independent of how $Z_{i}$ is spatially distributed in jammed packings.

### v2

@@ Introduction.
To demonstrate the instability of fragile states, two tests are carried out. First, we perform a compression-decompression perturbation, $\varphi\rightarrow\varphi+\delta\varphi\rightarrow\varphi$, where $\delta\varphi=10^{-8}$, with the energy minimized after each step. All fragile states ($43<Z<$ 4) become unjammed ($Z=0$) after this perturbation, while jammed states remain. In Fig. 4c, we plot the distribution $p^{\prime}(Z)$ after the compression-decompression perturbation. Independently, without any perturbation we simply regard all fragile states ($43<Z<$ 4) as unjammed ($Z=0$) and recalculate the distribution $\tilde{p}(Z)$. Figure 4c shows that the two distributions $p^{\prime}(Z)$ and $\tilde{p}(Z)$ perfectly coincide.
In the second test, we repeat CAQS simulations by systematically varying the threshold $f_{\rm th}$ in the criterion of the energy minimization algorithm. The fraction $F_{\rm F}$ of fragile states grows and $F_{\rm U}$ of unjammed states decays with increasing $f_{\rm th}$ (Fig. 4d), suggesting that many unjammed states become fragile under a looser force-balance condition $\overline{f}_{i}<f_{\rm th}$ with a larger $f_{\rm th}$. In contrast, the constant $F_{\rm J}$ shows that the definition of jammed states is insensitive to the algorithm parameter. The $p(Z)$ data with different $f_{\rm th}$ confirm this property: the fragile peak depends on $f_{\rm th}$ while the jammed peak is independent of $f_{\rm th}$ (see SI Fig. S1a). Note that in granular experiments Bi et al. (2011); Zhao et al. (2019), the friction may play the role of $f_{\rm th}$, i.e., the net inter-particle force could be balanced by the frictional force between particles and the supporting plate. According to this assumption, the probability of observing fragile states in experiments would depend on the particle-plate friction that can be changed by replacing the materials. It provides a protocol to examine our scenario of fragile states in future experimental studies.
@@
Figure 4: Stability tests. Average single-particle (a) potential energy $\overline{e}_{i}$ and (b) net force $\overline{f}_{i}$ as functions of $Z$. Lines in (a) and (b) are guides to the eye. (c) Comparison between $p^{\prime}(Z)$ obtained after the compression-decompression perturbation and $\tilde{p}(Z)$ with fragile states counted as unjammed states, for $N=1000$ at $\varphi=0.844$. (d) Fractions of four states as functions of $f_{\rm th}$. Data are obtained for $N=256$ systems at $\varphi=0.841$
, except for (c).
Figure 5: Distribution of the coordination number. (a) Probability distribution $p(Z)$ at a few different $\varphi$ for $N=1000$. (b) $p(Z)$ at $\varphi=0.843$ for a few different $N$. For better visualization, we do not show the unjammed delta-peak at $Z=0$ (see SI Fig. S1b for full distributions). (c) The fraction of jammed states $F_{\rm J}$ as a function of $\varphi$ for a few different $N$. The intersection of curves gives $\varphi_{\rm J}^{\infty}=0.8432$. (d) The data points of $F_{\rm J}$ with different $N$ collapse as a function of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$. The solid line represents fitting to Eq. (4), with two fitting parameters $u=0.041$ and $\sigma_{\varphi}=0.043$. The average coordination number $\langle Z\rangle$ is plotted as a function of $\varphi$ in (e) and of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$ in (f). The solid line in (f) represents $\langle Z\rangle=F_{\rm J}Z_{\rm J}^{*}$ using $u=0.041$, $\sigma_{\varphi}=0.043$ and $Z_{\rm J}^{*}=4.1$. Symbols in (b-f) have the same meanings.
Scaling analysis near the jamming transition.
[…]
(I) Standard first-order transitions without disorder. In this case, $F_{1}(x,N)$ has the scaling form $F_{1}(x,N)=\mathcal{F}_{1}[(x-x_{\rm c})N]$. The fluctuation-dissipation theorem ensures that the susceptibility can be equivalently computed from the response or from the fluctuations, i.e., $\chi_{\rm dis}=\chi_{\rm con}$. Examples in this category include the first-order transition in the $q$-state Potts model with $q>4$ where the control parameter is the temperature $T$, and that in the Ising model driven by an external magnetic field $h$ at temperatures below the critical temperature.
[…]
In Figs 5a and b, we plot $p(Z)$ for a few different $\varphi$ and $N$, showing that $p(Z)$ is always double-peaked across the jamming transition (see SI Fig. S1b), which is typical behavior of a first-order rather than a second-order transition. To analyze the scaling behavior of $p(Z)$, we consider a general form of first-order phase transitions Binder (1987a, b),
@@
where $\delta(Z)$ and $p_{\rm J}(Z)$ correspond to the unjammed and jammed peaks respectively. The asymptotic jamming density $\varphi_{\rm J}^{\rm\infty}=0.8432$ for $N\rightarrow\infty$ is determined by the intersection of $F_{\rm J}(\varphi)$ curves with different $N$ (see Fig. 5c). This value is consistent with the J-point density $\varphi_{\rm J}\approx 0.842-0.843$ given by previous studies O’hern et al. (2003); Vagberg et al. (2011), which is unambiguously below the maximum protocol-dependent jamming density $\varphi_{\rm j}\approx 0.8465$ reported in Ref. Kawasaki and Miyazaki (2020) and $\varphi_{\rm j}\approx 0.848$ obtained by CAQC in this study. The difference between $\varphi_{\rm J}^{\rm\infty}=0.8432$ by $\gamma_{\rm max}=0.7$ and $\varphi_{\rm J}^{\rm\infty}=0.8435$ by $\gamma_{\rm max}=1$ (see SI Fig. S2b and S3) is too small to conclude any systematic dependence on $\gamma_{\rm max}$. In Ref. Das et al. (2020), an unjamming pocket is reported in the phase diagram for $\gamma_{\rm max}<0.17$, while for larger $\gamma_{\rm max}$, $\varphi_{\rm J}$ is independent of $\gamma_{\rm max}$, consistent with our observations. These results also suggest that the jamming density obtained by cyclic shear is the lowest jamming density on the J-line of all possible jammed states Parisi and Zamponi (2010); Chaudhuri et al. (2010); Babu et al. (2021).
@@
We assume the fraction of jammed states $F_{\rm J}$ having the following scaling form, $F_{\rm J}(\varphi,N)=\mathcal{F}_{\rm J}\left[(\varphi-\varphi_{\rm J}^{\infty})N^{\lambda}\right]$. The value of the exponent $\lambda$ is important for understanding the nature of the transition. If the system were thermal, the fraction would be determined by the Boltzmann distribution, $F_{\rm J}\sim\exp\left(\frac{N\delta f}{k_{\rm B}T}\right)$, where $\delta f$ is the single-particle free energy difference between two phases Binder (1987b, a). Because the free energy is non-singular around a first-order phase transition, it can be expanded, giving $\delta f\sim(T-T_{\rm c})$ to the lowest order. Thus, if the jamming transition were a standard first-order phase transition, then $F_{\rm J}(\varphi,N)=\mathcal{F}_{\rm J}\left[(\varphi-\varphi_{\rm J}^{\infty})N\right]$, i.e., $\lambda=1$ (note that $\varphi$ is the control parameter of the jamming transition). However, our numerical data can not be collapsed using $\lambda=1$; in contrast, they can be perfectly collapsed using $\lambda=1/2$ (see Fig. 5d).
To understand the origin of $\lambda=1/2$ scaling, recall that the athermal jamming transition is not driven by the free energy difference between the two phases. Below we explain the scaling by the scenario of a first-order transition with quenched disorder. For a finite $N$, due to the presence of disorder in the packing structure, the jamming density $\varphi_{\rm J}^{N}$ should fluctuate around the asymptotic value $\varphi_{\rm J}^{\infty}$. Let us assume a simple Gaussian form of the distribution, $\rho(\varphi_{\rm J}^{N})\sim\exp\left[-\frac{(\delta\hat{\varphi}_{\rm J}+u)^{2}}{2\sigma_{\varphi}^{2}}\right],$ where $\delta\hat{\varphi}_{\rm J}=(\varphi_{\rm J}^{N}-\varphi_{\rm J}^{\infty})N^{1/2}$ follows the standard central limit theorem. Note that $\rho(\varphi_{\rm J}^{N})$ has been explicitly measured in the compression protocol O’Hern et al. (2002); O’hern et al. (2003): the width of $\rho(\varphi_{\rm J}^{N})$ scales as $w\sim N^{-0.55}$ in both 2D and 3D, supporting our assumption. The above assumption also predicts, $\varphi_{\rm J}^{\infty}-\varphi_{\rm J}^{N}\sim N^{-1/2}$, independent of the dimensionality. The jammed states are defined by $\varphi>\varphi_{\rm J}^{N}$, and thus
@@
where $\delta\hat{\varphi}=(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$. Equation (4) agrees well with the data in Fig. 5d. The dependence of $F_{\rm J}(\varphi,N)$ on shear parameters is examined in Fig. S2. The $F_{\rm J}(\varphi,N)$ data are independent of the maximum strain $\gamma_{\rm max}$, and weakly depend on the strain step $\delta\gamma$. Independent of $\delta\gamma$, the $F_{\rm J}(\varphi,N)$ data of different $N$ collapse when plotted as a function of the rescaled variable $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$ (see Figs. 5d and S2c for $\delta\gamma=10^{-2}$ and $10^{-1}$ respectively), showing robustness of the scaling function Eq. (4).
Figure 6: Susceptibilities. We plot (a) the disconnected susceptibility $\chi_{\rm dis}$ rescaled by $N$, (b) the connected susceptibility $\chi_{\rm con}$ rescaled by $N^{1/2}$, and (c) the ratio between $\chi_{\rm dis}$ and $\chi_{\rm con}^{2}$, as functions of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$. The solid black lines in (a-c) are Eqs. (5), (6), and the ratio between Eq. (5) and the square of Eq. (6), respectively. The dashed black line in (c) is the second-order expansion Eq. (7). The dashed red line in (c) is the constant term in Eq. (7), i.e., $\chi_{\rm dis}/\chi_{\rm con}^{2}=\pi\sigma_{\varphi}^{2}/2$. To draw these lines, we use $u=0.041$ and $\sigma_{\varphi}=0.043$ obtained from the fitting in Fig. 5d, and $Z_{\rm J}^{*}=4.1$ determined at $\varphi_{\rm J}^{\infty}=0.8432$ in Fig. 3d. No fitting is performed here. Symbols in (a-c) have the same meanings as indicated in the legend of (a). In (d) and (e), the susceptibilities of jammed states $\chi_{\rm J}$ are plotted as functions of $N$ for a few different $\varphi$, and are fitted to Eq. (8). The open squares in (d) are $\chi_{\rm J}$ data obtained in a small pressure window $0.00110.0009<P<$ 0.0011, corresponding to $\Delta Z\approx 0.15$ ($\varphi\approx 0.843-0.844$). The fitting parameter $\mu$ is plotted in (f) as a function of $\varphi$, where the error bar represents the fitting error.
Once the scaling behavior of $F_{\rm J}(\varphi,N)$ is obtained, it is easy to derive scaling forms of susceptibilities. From Eq. (3) and the definition of $\chi_{\rm dis}$, we obtain,
@@
From Eq. (3), one finds that $\langle Z\rangle\approx F_{\rm J}Z_{\rm J}^{*}$ and thus $\langle Z\rangle$ has the same scaling form as $F_{\rm J}$, i.e., $\langle Z\rangle(\varphi,N)=\mathcal{Z}[(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}]$, consistent with the data in Fig. 5e and 5f. Using $\chi_{\rm con}=Z_{\rm J}^{*}\frac{dF_{\rm J}}{d\varphi}$ and Eq. (4), we obtain,
@@
To the lowest order, Eq. (7) gives a scaling, $\chi_{\rm dis}\sim\chi_{\rm con}^{2},$ which is a key signature of the presence of quenched disorder. Comparing to the simulation data, Eq. (7) works well around the extreme point (see Fig. 6c). The agreement can be improved using higher-order expansions.
It seems that the non-Gaussian effect in the distribution $\rho(\varphi_{\rm J}^{N})$, which is neglected in the present analysis, is amplified in the data of $\chi_{\rm dis}/\chi_{\rm con}^{2}$, resulting in slight asymmetry. In SI Figs. S3-S6, we show that the scaling functions Eqs. (4)-(7) work in 3D, and for a different max strain $\gamma_{\max}=1$ in 2D. In 3D, we obtain $\varphi_{\rm J}^{\infty}=0.648$, consistent with the previously reported J-point density $\varphi_{\rm J}\approx 0.648$ of this model Das et al. (2020), which is the minimum jamming density on the J-line. For the same 3D model, the protocol-dependent jamming density is obtained up to $\varphi_{\rm j}=0.6616$ by the thermal annealing method Chaudhuri et al. (2010), and up to $\varphi_{\rm j}=0.661$ by the CAQS mechanical training method Das et al. (2020).
Next let us discuss the finite-size effects of the jammed peak $p_{\rm J}(Z)$ in Eq. (3). For standard first-order phase transitions, $\eta=1/2$ Binder (1987b, a). Thus in that scenario the susceptibility of the jamming peak would scale as, $\chi_{\rm J}\equiv N\sigma_{\rm J}^{2}=N\langle(Z-Z_{\rm J}^{*})^{2}\rangle_{\rm J}\sim{\rm constant}$, where $\sigma_{\rm J}^{2}$ is the variance of $p_{\rm J}(Z)$ and $\langle\ldots\rangle_{\rm J}$ represents the average restricted to the jammed states only. However, our $\chi_{\rm J}$ data disagree with this scaling (see Figs.  6d-f). At different $\varphi$, $\chi_{\rm J}$ can be fitted to a pow-law form $\chi_{\rm J}\sim\ell^{-\mu}$ Hexner et al. (2018, 2019), or
@@
where $N\sim\ell^{d}$. The exponent $\mu=d(2\eta-1)$ is a fitting parameter. At large or small $\varphi$ away from $\varphi_{\rm J}^{\rm\infty}$ (see Fig.  6f), the exponent is close to zero ($\mu\approx-0.2$), suggesting that the local coordination numbers $Z_{i}$ are uniformly distributed. The deviation from the uncorrelated behavior ($\mu=0$) is significant around $\varphi_{\rm J}^{\rm\infty}$, where $\mu\approx-0.8$ reaches the minimum.
@@
The result of $\mu$ shall be interpreted with care. In general, (i) $\mu=0$ corresponds to uniformity of $Z_{i}$, (ii) $\mu>0$ to hyperuniformity with a vanishing $\chi_{\rm J}$ in the thermodynamic limit, and (iii) $\mu<0$ to hyperfluctuations with a diverging $\chi_{\rm J}$ in the thermodynamic limit that typically appears at the critical point in a second-order phase transition. However, the negative $\mu$ in Fig. 6f is not due to the criticality of the jamming transition. Here it is essential to consider the volume fluctuations. The negative $\mu$ is obtained from a $\varphi$-controlled setup in our CAQS simulations. If we instead select configurations around a constant pressure $P$, then the fluctuations become significantly smaller, and $\mu\approx 0$ (Fig. 6d). This observation is consistent with a previous study Ikeda (2023), suggesting that the large fluctuation of $Z$ near $\varphi_{\rm J}^{\rm\infty}$ in the $\varphi$-controlled protocol might be originated from the fluctuation of $\varphi_{\rm J}^{N}$. In fact, careful measurement gives $\mu\simeq 1$ in a $P$-controlled protocol at jamming, confirming hyperuniformity Hexner et al. (2018, 2019). Ref. Hexner et al. (2019) also compares an ensemble of subsystems cut out from large packings, to an ensemble of whole systems with the same volume under periodic boundary conditions (similar to the $P$-controlled ensemble considered in this study): the contact fluctuations in the former are larger than those in the latter, and the convergence of these two are expected to occur in extremely large systems that are not attainable in current simulations – Due to this pre-asymptotic effect, the hyperuniform exponent $\mu\simeq 1$ is only observable in the first ensemble, while in the second ensemble, the apparent exponent is close to $\mu\approx 0$, consistent with our $P$-controlled data in Fig. 6d. In short, although the hyperuniform exponent $\mu=1$ is not directly observed, our results do not contradict the recently reported hyperuniform behavior of contact distributions at jamming.
We emphasize that the finite-size effects of the jammed peak $p_{\rm J}(Z)$ contribute negligibly to the scaling of $F(\varphi,N)$, $\chi_{\rm dis}(\varphi,N)$ and $\chi_{\rm con}(\varphi,N)$. In other words, the first-order nature of the jamming transition is independent of how $Z_{i}$ is spatially distributed in jammed packings.
In the scaling analysis presented above, $Z$ is treated as the essential order parameter. In principle, the same kind of scaling analysis can be applied to any physical quantity $A$ that varies discontinuously at jamming (e.g., $A$ can be the bulk modulus $B$ or the fraction of non-rattlers $f_{\rm NR}$). The distribution $p(A)$ should have the same form as Eq. (3), $p(A)=(1-F_{\rm J})\delta(A)+F_{\rm J}p_{\rm J}(A)$. With fragile states removed, generally $A=0$ below jamming because no contacts are formed (e.g., $B=0$ and $f_{\rm NR}=0$), and thus the unjammed peak is always a delta function. The fraction of jammed states $F_{\rm J}$ only depends on the constructed ensemble, which determines how the states are sampled, and thus its scaling form $F_{\rm J}(\varphi,N)=\mathcal{F}_{\rm J}[(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}]$ is independent of $A$. The jammed peak $p_{\rm J}(A)$ does depend on $A$, but as shown above, $p_{\rm J}(A)$ is irrelevant to the interested scaling behavior of $F_{\rm J},\chi_{\rm dis}$ and $\chi_{\rm con}$. Thus, our conclusion drawn from the scaling analysis should be robust and generic, independent of which parameter $A$ is chosen for the scaling analysis, as long as $A$ can reflect the discontinuous nature of jamming.